# FairGen: Controlling Sensitive Attributes for Fair Generations in Diffusion Models via Adaptive Latent Guidance

### v1

@@ 5.1 Bias Evaluation of FairGen
Table 2 demonstrates the bias scores $B$ (lower is better) and quality scores $Q$ (higher is better) for FairGen and the baselines on the HBE dataset on two types of text-to-image diffusion models, Stable Diffusion 2 (SD2) and Stable Diffusion XL (SDXL), across sensitive attributes gender, race, age, and their combination. Across all sensitive attributes and their combinations, FairGen consistently achieves the lowest bias scores, indicating its superior ability to mitigate multi-attribute bias without additional training. For instance, in case of gender, FairGen achieves a bias score of $0.231$ for SD2 and $0.267$ for SDXL, significantly outperforming all baselines. Similarly, when considering the combination of gender, race, and age, FairGen achieves the lowest bias scores of 0.601 on SD2 and 0.658 on SDXL. Notably, FairGen also sustains high generation quality in most of the cases compared to the baselines, with $Q$ scores that are competitive with or superior to vanilla generation (soft upper bound for quality scores without any interventions). These results underline FairGen’s ability to balance bias mitigation with image generation quality, especially in complex scenarios involving multiple intersecting sensitive attributes.
We also evaluate the effectiveness of FairGen and the baselines on the standard Stable Bias Luccioni et al. (2023) benchmark (for the occupation split and attribute “gender”) in Section E.1. First, we observe that FairGen outperforms all baselines both in bias reduction and quality preservation in Stable Bias. Second, we observe that in Stable Bias, the performance delta of the baselines with vanilla generation is significantly higher compared to their performance delta in HBE. For example, FairDiffusion method results in $34.5\%$ gender bias reduction in the Stable Bias benchmark, however, its gender bias reduction ($2.72\%$) is significantly lower in HBE where the prompts are more complex (Table 2). In contrast, FairGen achieves significant gender bias reduction in both Stable Bias ($79.95\%$) and HBE ($68.52\%$) datasets. This proves that our proposed FairGen is significantly more robust to prompt complexity compared to the baselines.
@@ 5.2 Effectiveness of FairGen with Different Target Generation Distributions
The results in Table 3 demonstrate that FairGen provides a robust and adaptable mechanism for controlling generation distributions to achieve targeted levels since the average bias is lower than other baselines at all levels. Specifically, FairGen demonstrates both the lowest average bias score and the smallest standard deviation, which indicates that it consistently maintains low bias across different target portions. This stability is critical, as it suggests that FairGen is not only effective at minimizing bias on average but also performs reliably across a wide range of scenarios. In contrast, while the finetune-based approach achieves relatively low bias scores among the baselines, its standard deviation is notably higher than that of FairGen. This higher variability implies that the finetune-based approach may be less predictable or stable when applied across different target portions. Methods like Vanilla generation and FairDiffusion also exhibit higher standard deviations, indicating a less consistent ability to manage bias across different target proportions.
@@ 5.3 FairGen with Different Diffusion Steps
In this part, we explore the impact of diffusion time steps to apply FairGen guidance on the effectiveness of bias mitigation and generation quality. The results in Table 4 demonstrate that applying latent guidance at the early diffusion stage (within the first $25\%$ time steps) does not effectively guide fair generations since later denoising downplays the early guidance, hence, it results in higher bias, however, with higher quality. Applying guidance at a later stage (last 25% time steps) degrades the alignment between generated images and input text resulting in lower quality. Therefore, we adopt guidance at the intermediate stage (middle 25% time steps) which ensures a desired balance between bias mitigation and generation quality.
@@ 5.4 Runtime Analysis and Other Ablations
We find that FairGen leads to marginal runtime overhead compared to the baselines while reducing the bias significantly (Section E.3). We perform an ablation for the SFT and DPO steps for training the attribute-aware generator $L$ (Appendix E.2) and find that SFT+DPO version performs the best. We provide visualization examples in Section E.4.
@@ 6 Related Work
Bias Evaluation in Diffusion Models. Evaluation of bias in text-to-image diffusion models has gained much interest recently. Numerous works have studied demographic biases (e.g., gender, race, etc.) in different domains such as occupation, physical characteristics, and so on (Bakr et al., 2023; Lee et al., 2024; Cui et al., 2023; Wan and Chang, 2024; Wan et al., 2024; Luccioni et al., 2023; Naik and Nushi, 2023). These studies focus on constructing attributed prompts (e.g., photo of a <objective>) to probe the text-to-image models for any bias towards a specific attribute value (e.g., towards “male” when generating images of engineers). However, current studies overlook many domains such as healthcare, finance, and everyday activities and they rely on simplistic prompts for probing the models, hence, fail to capture the complexity and nuance of real-world user inputs. We address the two limitations and propose the HBE benchmark which covers a broader range of sensitive attributes and domains, sampled from realistic statistical distribution of user prompts and rigorously filtered.
@@ 7 Conclusion
FairGen presents a significant advancement in addressing the challenge of generative bias in diffusion models. By integrating adaptive latent guidance with a global memory, FairGen not only mitigates bias but also maintains high-quality image generation. The dynamic adjustment of latent space attributes and the use of generation statistics make it controllable in multi-attribute scenarios and adaptive to different target distributions. Extensive evaluations and ablations demonstrate FairGen’s superior performance in reducing bias compared to state of the art techniques, while also offering precise and flexible control over targeted distributions. This work paves the way for more socially responsible applications of diffusion models in real-world use cases.
@@ Appendix B Preliminaries
where $f({\mathbf{x}},t):{\mathbb{R}}^{n}\mapsto{\mathbb{R}}^{n}$ is the drift coefficient characterizing the shift of the distribution, $g(t)$ is the diffusion coefficient controlling the noise scales, and $\mathbf{w}$ is the standard Wiener process. The reverse process is characterized via the reverse time SDE of Equation 6:
@@
Since the SDE formulation in Equation 6 is typically discretized for numerical computations, we basically consider the discrete process formulation.
@@ Appendix C Efficacy of Model Judge
Across the evaluations, we detect the attributes of persons in generated images by performing question-answering with the InstructBLIP-2 model. Specifically, we directly ask the vision-language model to do a classification task for gender, race, or age. In this part, we evaluate the alignment between the model judge and the human judge. Using Amazon Sagemaker GroundTruth platform, we invited Amazon Mechnical Turk workers to annotate the gender, race, and age for $100$ images. For each of the $100$ images, we obtained the labels across different sensitive attributes. We then computed the efficacy of model judge in Table 6. The results show that model judge by InstructBLIP-2 shows overall desirable attribute detection performance. The human evaluators were compensated according to California minimum hourly pay rate. Demographically, they were English speaking and 18+ in age. They were warned that the generations are from a machine learning model and may contain sensitive or unexpected components.
@@ E.1 Evaluation on the Stable Bias Dataset
Table 7 presents the bias score $B$ and quality score $Q$ for various bias mitigation methods evaluated on the stable bias occupation dataset (Luccioni et al., 2023) for the attribute “gender”. Among all methods, FairGen demonstrates the most significant bias reduction, achieving a bias score of 0.160, which is substantially lower than the other baselines. This again indicates its superior performance in mitigating bias in various datasets. While the fine-tune-based method also shows notable bias reduction with a score of 0.392, FairGen surpasses it by a large margin and is also more flexible to the change of target portions. Additionally, FairGen maintains a high generation quality score ($Q=0.297$), which is competitive with vanilla generation ($Q=0.303$) and higher than most other approaches. This indicates that FairGen strikes an effective balance between minimizing bias and preserving image quality.
Table 7: Bias score $B$ ($\downarrow$) and quality score $Q$ ($\uparrow$) on stable bias occupation dataset for the attribute “gender”.
Method $B$ $Q$
Vanilla generation 0.798 0.303
Prompt intervention 0.637 0.267
Fine-tune-based method 0.392 0.281
FairDiffusion 0.523 0.284
FairGen 0.160 0.297
@@ E.2 Effectiveness of SFT and DPO
During the training of guidance prompt generation model in Section 2.2, we leverage a dual-phase mechanism: SFT which imposes attribute-aware prompt generation and DPO which further refines model with fairness generation utility feedback. In this part, we directly verify the effectiveness of SFT and DPO. We prompt LLM to add attribute specification as a baseline and compare it with FairGen (SFT) and FairGen (SFT+DPO). As shown in Table 8, the baseline LLM prompting achieves a bias score $B$ of 0.203 and a quality score $Q$ of 0.298. When SFT is applied, we observe a reduction in bias to 0.168 while maintaining a similar quality score of 0.299, indicating that SFT benefits LLM capacity for attribute-aware guidance prompt generation. Furthermore, adding DPO to SFT further reduces the bias score to 0.160, while keeping the fairness quality virtually unchanged, suggesting that DPO enhances the model by including additional feedback on quality of guidance prompts, which benefits the model to capture more nuanced correlations between prompt structures and fairness utilities.
@@ E.3 Runtime Analysis
We also evaluate the runtime of FairGen and other bias mitigation baselines in both the training phase and inference phase in Table 9. As a training-free method, FairGen induces no training computational costs. In the inference stage, although FairGen induces $1+2|\mathcal{A}|$ noises estimates in each diffusion step, where $|\mathcal{A}|$ is the number of sensitive attributes, the adaptive guidance is only enforced at a small portion of intermediate diffusion steps (details in Section 5.3). Additionally, the noise estimates for different attributes are independent and parallelized in the inference. Therefore, FairGen only leads to marginal runtime overhead compared to the baselines while mitigating the bias significantly.
Table 9: Comparison of runtime (hours) between FairGen and other bias mitigation baselines on stable diffusion 2 model on HBE benchmark. instance wise runtime
@@ E.4 Visualization Examples
In Figure 2, we present a series of image generations produced by FairGen, demonstrating its ability to precisely control the gender attribute while maintaining a high level of image fidelity. The figure highlights several key aspects of our model’s capabilities: (1) FairGen effectively adjusts the gender attribute across all generations, ensuring a balanced distribution between male and female representations. (2) The generated images exhibit high fidelity, preserving fine details in both the subjects and their surrounding environment. This demonstrates the robustness of FairGen in generating photorealistic images, even under conditions where specific attributes (e.g., gender) are modified. (3) Importantly, FairGen is able to control gender attributes without intervening with the background elements or scene composition.
Figure 2: Image generations by FairGen to control a balanced gender distribution.

### v2

@@ 5.1 Bias Evaluation of FairGen
Table˜2 demonstrates the bias scores $B$ (lower is better) and quality scores $Q$ (higher is better) for FairGen and the baselines on the HBE dataset on three types of text-to-image diffusion models, Stable Diffusion 2 (SD2), Stable Diffusion XL (SDXL), and Stable Diffusion 3.5 Large (SD-3.5-large), across sensitive attributes gender, race, age, and their combination. Across all sensitive attributes and their combinations, FairGen consistently achieves the lowest bias scores, indicating its superior ability to mitigate multi-attribute bias without additional training. For instance, in case of gender, FairGen achieves a bias score of $0.231$ for SD2 ,$0.267$ for SDXL, and $0.118$ for SD-3.5-large, significantly outperforming all baselines. Similarly, when considering the combination of gender, race, and age, FairGen achieves the lowest bias scores of 0.601 on SD2, 0.658 on SDXL, and 0.397 on SD-3.5-large. Notably, FairGen also sustains high generation quality in most cases compared to the baselines, with $Q$ scores that are competitive with or superior to vanilla generation (soft upper bound for quality scores without any interventions). These results underline FairGen’s ability to balance bias mitigation with image generation quality, especially in complex scenarios involving multiple intersecting sensitive attributes.
We also evaluate the effectiveness of FairGen and the baselines on the standard Stable Bias Luccioni et al. (2023) benchmark (for the occupation split and attribute “gender”, “race", and “age") in Section˜E.1. We also observe that FairGen achieves significant gender bias reduction and better generation quality compared to the baselines.
@@ 5.2 Effectiveness of FairGen with Different Target Generation Distributions
The results in Table˜3 demonstrate that FairGen provides a robust and adaptable mechanism for controlling generation distributions to achieve targeted levels since the average bias is lower than other baselines at all levels. Specifically, FairGen demonstrates both the lowest average bias score and the smallest standard deviation, which indicates that it consistently maintains low bias across different target portions. This stability is critical, as it suggests that FairGen is not only effective at minimizing bias on average but also performs reliably across a wide range of scenarios. In contrast, while the finetune-based approach achieves relatively low bias scores among the baselines, its standard deviation is notably higher than that of FairGen. This higher variability implies that the finetune-based approach may be less predictable or stable when applied across different target portions. Methods like Vanilla generation and FairDiffusion also exhibit higher standard deviations, indicating a less consistent ability to manage bias across different target proportions.
@@ 5.3 FairGen with Different Diffusion Steps
In this part, we explore the impact of diffusion time steps to apply FairGen guidance on the effectiveness of bias mitigation and generation quality. The results in Table˜4 demonstrate that applying latent guidance at the early diffusion stage (within the first $25\%$ time steps) does not effectively guide fair generations since later denoising downplays the early guidance, hence, it results in higher bias, however, with higher quality. Applying guidance at a later stage (the last 25% of time steps) degrades the alignment between the generated images and the input text, which results in lower quality. Therefore, we adopt guidance at the intermediate stage (middle 25% time steps) which ensures a desired balance between bias mitigation and generation quality.
@@ 5.4 Runtime Analysis and Other Ablations
We report the runtime of FairGen and other baselines in Table˜9 (Section˜E.3). Since FairGen is training-free, it incurs no additional training cost. During inference, although it introduces extra noise estimates at each diffusion step, adaptive guidance is applied during only a small subset of intermediate steps (Section˜5.3). These estimates are attribute-independent and parallelizable, resulting in only a marginal runtime overhead while achieving significant bias reduction.
We ablate the SFT and DPO steps for training the attribute-aware generator $L$ (Section˜E.2) and find that combining both yields the best performance. Visualization examples are shown in Section˜E.4.
@@ 6 Related Work
Bias Evaluation in Diffusion Models. Evaluation of bias in text-to-image diffusion models has gained significant interest recently. Numerous works have studied demographic biases in different domains such as occupation, physical characteristics, and so on (Bakr et al., 2023; Lee et al., 2024; Cui et al., 2023; Wan and Chang, 2024; Wan et al., 2024; Luccioni et al., 2023; Naik and Nushi, 2023). These studies focus on constructing attributed prompts (e.g., photo of a <objective>) to probe the text-to-image models for any bias towards a specific attribute value (e.g., towards “male” when generating images of engineers). However, current studies overlook many domains such as healthcare, finance, and everyday activities and they rely on simplistic prompts for probing the models, hence, fail to capture the complexity and nuance of real-world user inputs. We address the two limitations and propose the HBE benchmark which covers a broader range of sensitive attributes and domains, sampled from realistic statistical distribution of user prompts and rigorously filtered.
@@ 7 Conclusion
FairGen introduces a substantial improvement in mitigating generative bias in diffusion models. By integrating adaptive latent guidance with a global memory, it effectively reduces bias while preserving high-quality image generation. The dynamic adjustment of latent attributes and use of generation statistics enable precise control in multi-attribute settings and adaptability to varying target distributions. Extensive evaluations and ablations show that FairGen consistently outperforms existing approaches in both bias mitigation and controllability, offering a practical step toward more socially responsible diffusion-based applications.
@@
Computation overhead. While FairGen introduces relatively modest runtime overhead (shown in Table˜9), we acknowledge that it may still affect scenarios where throughput is critical. It would be interesting future work to further mitigate the tradeoff between the fairness of diffusion generation distribution and the inference efficiency.
@@ Appendix B Preliminaries
where $f({\mathbf{x}},t):{\mathbb{R}}^{n}\mapsto{\mathbb{R}}^{n}$ is the drift coefficient characterizing the shift of the distribution, $g(t)$ is the diffusion coefficient controlling the noise scales, and $\mathbf{w}$ is the standard Wiener process. The reverse process is characterized via the reverse time SDE of Equation˜6:
@@
Since the SDE formulation in Equation˜6 is typically discretized for numerical computations, we basically consider the discrete process formulation.
@@ Appendix C Efficacy of Model Judge
Across the evaluations, we detect the attributes of persons in generated images by performing question-answering with the InstructBLIP-2 model. Specifically, we directly ask the vision-language model to do a classification task for gender, race, or age. In this part, we evaluate the alignment between the model judge and the human judge. Using Amazon Sagemaker GroundTruth platform, we invited Amazon Mechnical Turk workers to annotate the gender, race, and age for $100$ images. For each of the $100$ images, we obtained the labels across different sensitive attributes. We then computed the efficacy of model judge in Table˜6. The results show that model judge by InstructBLIP-2 shows overall desirable attribute detection performance. The human evaluators were compensated according to California minimum hourly pay rate. Demographically, they were English speaking and 18+ in age. They were warned that the generations are from a machine learning model and may contain sensitive or unexpected components.
@@ E.1 Evaluation on the Stable Bias Dataset
Table˜7 presents the bias score $B$ and quality score $Q$ for various bias mitigation methods evaluated on the Stable Bias occupation dataset (Luccioni et al., 2023) for the different attributes. Among all methods, FairGen demonstrates the most significant bias reduction, achieving bias scores substantially lower than the other baselines. This again indicates its superior performance in mitigating bias in various datasets. While the fine-tune-based method also shows notable bias reduction with a score, FairGen surpasses it by a large margin and is also more flexible to the change of target portions. Additionally, FairGen maintains a high generation quality score, which is competitive with vanilla generation and higher than most other approaches. This indicates that FairGen strikes an effective balance between minimizing bias and preserving image quality.
Table 7: Bias score $B$ ($\downarrow$) and quality score $Q$ ($\uparrow$) on the Stable Bias occupation dataset for attributes gender, age, and race.
Method Gender Age Race
$B$ $Q$ $B$ $Q$ $B$ $Q$
Vanilla generation 0.798 0.303 0.925 0.327 0.839 0.285
Prompt intervention 0.637 0.267 0.782 0.268 0.583 0.246
Fine-tune-based method 0.392 0.281 0.374 0.260 0.240 0.251
FairDiffusion 0.523 0.284 0.342 0.258 0.320 0.248
FairGen (FairGen) 0.160 0.297 0.206 0.307 0.189 0.283
@@ E.2 Effectiveness of SFT and DPO
During the training of guidance prompt generation model in Section˜2.2, we leverage a dual-phase mechanism: SFT which imposes attribute-aware prompt generation and DPO which further refines model with fairness generation utility feedback. In this part, we directly verify the effectiveness of SFT and DPO. We prompt LLM to add attribute specification as a baseline and compare it with FairGen (SFT) and FairGen (SFT+DPO). As shown in Table˜8, the baseline LLM prompting achieves a bias score $B$ of 0.203 and a quality score $Q$ of 0.298. When SFT is applied, we observe a reduction in bias to 0.168 while maintaining a similar quality score of 0.299, indicating that SFT benefits LLM capacity for attribute-aware guidance prompt generation. Furthermore, adding DPO to SFT further reduces the bias score to 0.160, while keeping the fairness quality virtually unchanged, suggesting that DPO enhances the model by including additional feedback on quality of guidance prompts, which benefits the model to capture more nuanced correlations between prompt structures and fairness utilities.
@@ E.3 Runtime Analysis
We also evaluate the runtime of FairGen and other bias mitigation baselines in both the training phase and inference phase in Table˜9. As a training-free method, FairGen induces no training computational costs. In the inference stage, although FairGen induces $1+2|\mathcal{A}|$ noises estimates in each diffusion step, where $|\mathcal{A}|$ is the number of sensitive attributes, the adaptive guidance is only enforced at a small portion of intermediate diffusion steps (details in Section˜5.3). Additionally, the noise estimates for different attributes are independent and parallelized in the inference. Therefore, FairGen only leads to marginal runtime overhead compared to the baselines while mitigating the bias significantly.
Table 9: Comparison of runtime (hours) between FairGen and other bias mitigation baselines on stable diffusion 2 model on HBE benchmark.
@@ E.4 Visualization Examples
In Figure˜2, we present a series of image generations produced by FairGen, demonstrating its ability to precisely control the gender attribute while maintaining a high level of image fidelity. The figure highlights several key aspects of our model’s capabilities: (1) FairGen effectively adjusts the gender attribute across all generations, ensuring a balanced distribution between male and female representations. (2) The generated images exhibit high fidelity, preserving fine details in both the subjects and their surrounding environment. This demonstrates the robustness of FairGen in generating photorealistic images, even under conditions where specific attributes (e.g., gender) are modified. (3) Importantly, FairGen is able to control gender attributes without intervening with the background elements or scene composition.
Figure 2: Image generations by FairGen to control a balanced gender distribution by SD2.
Figure 3: Image generations by FairGen to control a balanced gender distribution by SD-3.5-Large.